### v2

@@ 5 Conclusions
In this research, we have demonstrated the utility of a hybrid quantum-classical approach for a close-to-real-life AGVs scheduling problem that can be implemented in a real factory environment. We have not yet demonstrated quantum advantage, but the results of one of the hybrid quantum-classical approaches are close to those of CPLEX, which is widely recognized as the benchmark classical solver. As such hybrid approaches are expected to improve in the near future, we may be on the edge of quantum advantage. While larger problems of hundreds of AGVs that are expected to arise from the developments of the concept of industry $4.0$ (yielding several thousands of variables in our model) may remain out of the range of classical algorithms, future hybrid quantum-classical algorithms can be expected to have the potential to cope with them.
The CQM solver even for large problems provides the spectrum of distinguishable solutions, that can be the input to multi cases DSS system or sophisticated stochastic scheduling approach. Besides, CPLEX (with limited time) is faster and closer to optimality it provides only one solution. The experience with BQM underlines the importance of systematic penalty determination [33] [32].
@@ Acknowledgements
Z.P., B.G., S.D., K.D. - conceptualization, K.D., T.Ś. - preparing experiments, T.Ś. - running experiments, K.D., Z.P., Ł.P., M.K., B.G., S.D. - data analysis, K.D., M.K, T.Ś. - manuscript writing, Z.P., Ł.P., B.G., S.D. - manuscript supervision. All authors reviewed the manuscript.
Appendix A QUBO model
The most general form of inequalities in our model is in Eq. (15) or Eq. (16). Based on these, we present detailed path of transformation of these inequalities into QUBO. Let us contract $t$-variables and $y$-variables to vectors with elements $t_{j}$ and $y_{i}$ (the same with constant terms $c_{i}$). Then, from the aforementioned equations (or similar), we have:
$t_{j}-t_{j^{\prime}}-My_{i}+\xi_{i}=-c_{i}.$ (29)
To replace inequalities with equalities, we use slack variables
as $\min(t_{j})=v_{j}$, and $\max(t_{j})=v_{j}+d_{\max}$ from Eq. (5), and we expect $\upsilon_{j^{\prime}}=\upsilon_{j}+c_{i}$. Observe that some of the inequalities such as these of the minimal passing time constraints in Eq. (14) do not have order variables ($y$ or $z$), for such variables $\bar{\xi}_{i}=d_{\max}$.
To convert the constrained problem in Eq. (29) into the unconstrained one, we use the penalty method: each constraint violation is penalized by the hard constraint penalty $p>0$. Such a penalty has to be sufficiently large not to be overruled by the objective. Then, the optimization problem has the form:
The size of the problem (in terms of the number of QUBO variables) scales linearly with the number of zones and quadratically in the number of AGVs. The number of constraints per variable (the mean vertex degree of the graph model) is tied to the expansion of quadratic terms in Eq. (31). In Tab. 3, sizes of actual problems in terms of Ising variables (derived directly from QUBO) are presented.
For the $15$ AGVs and $7$ zones problem we have $696$ variables (upper limit according to Eq. (12) was $945$), and $1378$ linear constraints (upper limit according to Eq. (28) (27) was $5513$). These limits are quadratic in the number of AGVs and linear in the number of zones. We may expect the problem size to be quadratic with the number of AGVs as the number of zones is constant. If, however, the number of AGVs increases further, more zones will appear where collisions occur, and therefore, the problem size will scale worse than quadratic.
$|J|$ / $|S|$
n.o. qubits
n.o. quadratic couplings
edges
edge density [41] $\frac{edges}{fullgraph}$ n.o. linear fields
2 / 4 122 1066 0.14 121
4 / 4 268 2644 0.07 267
6 / 7 796 11954 0.04 782
7 / 7 1204 19084 0.03 1183
12 / 7 6357 116422 0.006 6250
15 / 7 7343 134415 0.005 7219
Table 3: Sizes of analyzed AGV problems, in terms of Ising approach. Observe, that for the density is of the order of magnitudes lower, than the one for corresponding full graph. Concerning the largest problem, the factor is $0.005$.
Appendix B Details of one particular AGV problem
@@
In practice, the above will be done by step $2$ of Algorithm in Section 3.1. In objective function in Eq. (13) for this computational example each AGV is assigned an equal weight of $w_{j}=1$. Further, we set $d_{\max}=40$.
The most important step of the Algorithm in Section 3.1 is step $5$: solving the optimization problem. The CQM hybrid solver yields many solutions at a time. The best solution conicides with the optimal one also provided by CPLEX. However other solutions can also be of use for the decision support system. The soltuions are presented in the form of a simplified time-space diagram (in which known points are connected with lines) in Fig. 7.
Refer to captionRefer to captionRefer to caption

### v3

@@ 5 Conclusions
In this work, we have demonstrated the utility of a hybrid quantum-classical approach for a near-real-life AGVs scheduling problem that can be implemented in a real factory environment. While we have not yet demonstrated a quantum advantage, one of the hybrid quantum-classical approaches produced results close to those of CPLEX, which is widely recognized as the benchmark classical solver. As these hybrid approaches are expected to improve in the near future, we may be on the brink of, at least, achieving quantum advantage.
Larger problems involving hundreds of AGVs, anticipated with the development of Industry 4.0 (resulting in several thousand variables in our model), may remain beyond the capabilities of classical algorithms. However, future hybrid quantum-classical algorithms are expected to have the potential to handle these larger problems.
The CQM solver, even for large problems, provides a range of diverse solutions, which can be used as input for multi-case decision support systems or sophisticated stochastic scheduling approaches. In contrast, while CPLEX (with limited time) is faster and closer to optimality, it provides only one solution. Our experience with BQM highlights the importance of systematic penalty determination as discussed in recent studies [36] [35].
We recognize the significant methodological value in contributions that utilize open-source tools or develop custom hybrid classical-quantum workflows. To our knowledge, no production-quality general hybrid quantum-classical solvers that input native ILP are currently available. In contrast, for the QUBO implementation, we applied the open-source SBM solver, which outperforms the BQM solver in terms of solution quality but still fails to return feasible solutions (except for the smallest cases). Since the SBM solver is quantum-inspired and does not use quantum resources, our findings suggest that quantum annealing does not play a predominant role in the iterative hybrid process, aligning with the "Quantum Backstops" hypothesis. These observations highlight the need for more transparent hybrid quantum-classical or quantum-inspired solvers to effectively benchmark quantum (or quantum-inspired) computation. We believe it is important for the quantum computing scientific community to systematically work on these developments.
@@ Acknowledgements
Z.P., B.G., S.D., K.D. - conceptualization, K.D., T.Ś., J.P., A.P. - preparing experiments, T.Ś., J.P., A.P. - running experiments, K.D., Z.P., Ł.P., M.K., J.P., A.P., B.G., S.D. - data analysis, K.D., M.K, T.Ś., J.P. - manuscript writing, Z.P., Ł.P., B.G., S.D. - manuscript supervision. All authors reviewed the manuscript.
Appendix A Details of one particular AGVs problem
@@
In practice, the above will be done by step $2$ of Algorithm in Section 3.1. In the objective function in Eq. (13) for this computational example each AGV is assigned an equal weight of $w_{j}=1$. Further, we set $d_{\max}=40$.
The most important step of the Algorithm in Section 3.1 is step $5$: solving the optimization problem. The CQM hybrid solver yields many solutions at a time. The best solution coincides with the optimal one also provided by CPLEX. However other solutions can also be of use for the decision support system. The solutions are presented in the form of a simplified time-space diagram (in which known points are connected with lines) in Fig. 7.